You conduct a trial where 75 people are randomly selected and asked if they prefer coffee from MD. Your responses indicate that 60% of them did prefer coffee from MD.

Find the mean and standard deviation.
Answer #1

Solution :

Given that,

Using binomial distribution,

Mean = = n * p = 75 * 0.60 = 45

Standard deviation = = n * p * q = 75 * 0.60 * 0.40 = 18 = 4.2426